# Nonlinear trident using WKB and worldline instantons

### v1

@@ V Gamow-Sommerfeld suppression
$P(\text{with Coulomb})=C^{2}P(\text{without Coulomb})$ (203)
@@
Our focus in this section will be on the repulsive case, $q_{1}q_{1}>0$, but we first mention that the attractive case, $q_{1}q_{2}<0$, has an interesting $v\to 0$ limit. Consider annihilation of an $e^{\scriptscriptstyle+}e^{\scriptscriptstyle-}$ pair into $\mu^{\scriptscriptstyle+}\mu^{\scriptscriptstyle-}$, $\tau^{\scriptscriptstyle+}\tau^{\scriptscriptstyle-}$, $p\bar{p}$ or some other pair of particles. At the threshold, $v\to 0$ in the final state and
$P(\text{without Coulomb})\propto v\to 0\;,$ (205)
@@
so $P(\text{with Coulomb})$ remains nonzero as $v\to 0$, i.e. as a function of the initial energy the cross section has a discontinuous step at threshold [34, 35, 36, 37]. Such steps have been observed in experiments [38, 39, 40] (see also references in [40], e.g. for discussions about the existence of steps even for neutral baryon pairs). Sommerfeld enhancement has also been important in studies of dark matter, see [41] for a pedagogical explanation and for further references.
@@
$S_{j}=\int\mathrm{d}t\,m\sqrt{1-v_{j}^{2}}\approx\int\mathrm{d}t\left(-\frac{mv_{j}^{2}}{2}\right)+\text{const.}$. (212)

### v2

@@ V Gamow-Sommerfeld suppression
[…]
@@
Our focus in this section will be on the repulsive case, $q_{1}q_{1}>0$, but we first mention that the attractive case, $q_{1}q_{2}<0$, has an interesting $v\to 0$ limit. Consider the annihilation of an $e^{\scriptscriptstyle+}e^{\scriptscriptstyle-}$ pair into $\mu^{\scriptscriptstyle+}\mu^{\scriptscriptstyle-}$, $\tau^{\scriptscriptstyle+}\tau^{\scriptscriptstyle-}$, $p\bar{p}$ or some other pair of particles. At the threshold, $v\to 0$ in the final state and
[…]
@@
so P(with Coulomb)P($\text${with Coulomb}) remains nonzero as $v\to 0$, i.e. as a function of the initial energy the cross section has a discontinuous step at threshold [37, 38, 39, 40]. Such steps have been observed in experiments [41, 42, 43] (see also references in [43], e.g. for discussions about the existence of steps even for neutral baryon pairs). Sommerfeld enhancement has also been important in studies of dark matter, see [44] for a pedagogical explanation and for further references.
@@
$S_{j}=\int\mathrm{d}t\,m\sqrt{1-v_{j}^{2}}\approx\int\mathrm{d}t\left(-\frac{mv_{j}^{2}}{2}\right)+\text{const.}$ (269)
@@ Appendix A AAMLR
While our final result (266) agrees with Eq. (2.29) in [46], we cannot make a concrete comparison with their implicit regularization or their indirect discussion about renormalization. However, from the perturbative treatment in the appendix of [46], it seems that they might have had a more traditional perturbative renormalization in mind.